OpenStudy (anonymous):

how much does a salesman get if he was able to sell P 2,500,000.00 worth of property with 8%? can i get some solutions?

OpenStudy (tkhunny):

It is hoped that there is only one solution. What would you get if it were 100% Commission?

OpenStudy (anonymous):

yes sir

OpenStudy (tkhunny):

You didn't answer my question.

OpenStudy (anonymous):

sir this is my homework today.. it is 100% comission...

OpenStudy (anonymous):

???

OpenStudy (tkhunny):

100% Commission - 2,500,000 * 1.00 = 2,500,000 50% Commission - 2,500,000 * 0.50 = 1,250,000 25% Commission - 2,500,000 * 0.15 = 625,000 10% Commission - 2,500,000 * 0.10 = 250,000 Are we seeing a pattern?